# Modeling of AR 12760 with GX Simulator and Evidence for the Extended Transition Region in Peripheral Active Region Loops

### v1

@@ 5 Summary and Conclusions
We used GX Simulator to model the EUV flux from a small, quiescent active region averaged over a 4 hour period in order to determine the dependence of the volumetric heating, $\langle Q\rangle$, on the average magnetic field strength along the loop, $B_{\mbox{\scriptsize avg}}$, and loop length, $L$, assuming $\langle Q\rangle\propto B_{\mbox{\scriptsize avg}}^{a}L^{-b}$. We found that for the AIA bands dominated by emission with $\log T\mbox{(K)}\gtrsim 6.3$ (211 and 335 Å) the larger scale intensity distribution was reasonably well fit, although the detailed structure was different due to difficulties in the magnetic modeling of an active region with low intensity field. Using a relatively low resolution version of the model (8″) we found the best fit for 211 Å has $\langle Q\rangle\approx 7\times 10^{-3}(B_{\mbox{\scriptsize avg}}/{100\mbox{G}})^{1.5}(L/{10^{9}\mbox{cm}})^{-1}$ erg cm-3 s-1, and for 335 Å the best fit is $\langle Q\rangle\approx 4\times 10^{-3}(B_{\mbox{\scriptsize avg}}/{100\mbox{G}})^{0.5}(L/{10^{9}\mbox{cm}})^{-1.5}$ erg cm-3 s-1. However, for all fits there is a range of parameter values that give qualitatively reasonable fits lying roughly in a line in ($a,b$) space, indicative of a dependence between $L$ and $B_{\mbox{\scriptsize avg}}$. This line was approximately defined by $b\approx 1.6-0.6a$. This relationship is consistent with a $B_{\mbox{\scriptsize avg}}$ dependence on $L$.
@@
We also found that for EUV lines showing a cooler (but still optically thin) plasma the models did not replicate the emission from loops along the edge of the active region, which dominates the emission in 171 and 131 Å bands and also contributes significantly to the 94 and 193 Å bands. We conclude that this emission is due to relatively high-altitude transition region emission in these long loops. EBTEL defines the transition region to be the section of the loop in which thermal conduction serves to heat the plasma rather than cool it (as occurs in the coronal section), and as such the transition region generally fills a length about 5% of the full loop length from the footpoint. GX Simulator and the EBTEL model informing it assume that all transition region emission comes from one voxel at the footpoint, which is not sufficient to model the spatial distribution of the emission in these AIA bands. Thus we conclude that it is important in modeling active regions to consider the spatial extent of the upper transition region as a function of temperature. The cooler part of the transition region is confined to low altitudes close to the footpoints.
@@
T.A.K. and J.A.K. were supported by the GSFC Heliophysics Internal Scientist Funding Model (competitive work package program). This work is also part of the effort of the Committee on Space Research International Space Weather Actions Team (COSPAR-ISWAT) S2-06: Origins of the Spectral Irradiance and its Intermediate Timescale Variability. GX Simulator is a package available as part of Solar Soft (Freeland & Handy, 1998) and also on github at https://github.com/Gelu-Nita/GX_SIMULATOR. G.D.F. and G.M.N. were partly supported by the NSF grants RISE-2324724 and AST-2206424 and NASA grant 80NSSC23K0090 to the New Jersey Institute of Technology.
Appendix A AIA Model Grids
@@
This appendix contains more examples of calculated models that give the lowest $\sum\delta^{2}$ values as a function of $q_{0}$ for each of a grid of ($a,b$) values.
Figure 17: Thumbnails of model 211 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for calculations using Magnetic Model 1, a model/data comparison resolution of 2″, and including the corona only. The actual 211 Å image at the same resolution is at the bottom left. The $a=2$, $b=0.5$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. Maps with $\sum\delta^{2}$  within 10% and 20% of the minimum have titles and borders in orange and red respectively.
@@
Figure 18: Thumbnails of model 211 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for calculations using Magnetic Model 1, a model/data comparison resolution of 8″, and including both the corona and transitions region. The actual 211 Å image at the same resolution is on the bottom left. The $a=1.5$, $b=1$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. Maps with $\sum\delta^{2}$  within 10% and 20% of the minimum have titles and borders in orange and red respectively.
@@
Figure 19: Thumbnails of model 171 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for calculations using Magnetic Model 2, a model/data comparison resolution of 2″, and including the corona only. The actual 171 Å image at the same resolution is at the bottom left. The $a=1.0$, $b=0$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. Maps with $\sum\delta^{2}$  within 10% and 20% of the minimum have titles and borders in orange and red respectively.
@@
Figure 20: Thumbnails of model 211 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for calculations using Magnetic Model 1, a model/data comparison resolution of 8″, and including both the corona and transitions region. The actual 335 Å image at the same resolution is on the bottom left. The $a=0.5$, $b=1.5$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. Maps with $\sum\delta^{2}$  within 10% and 20% of the minimum have titles and borders in orange and red respectively.

### v2

@@ V Summary and Conclusions
We used GX Simulator to model the EUV flux from a small, quiescent active region averaged over a 4 hour period in order to determine the dependence of the volumetric heating, $\langle Q\rangle$, on the average magnetic field strength along the loop, $B_{\mbox{\scriptsize avg}}$, and loop length, $L$, assuming $\langle Q\rangle\propto B_{\mbox{\scriptsize avg}}^{a}L^{-b}$. We found that for the AIA bands dominated by emission with $\log T\mbox{(K)}\gtrsim 6.3$ (211 and 335 Å) the larger spatial scale intensity distribution was reasonably well fit, although the detailed structure was different due to difficulties in the magnetic modeling of an active region with relatively weak field. Using a relatively low-resolution version of the model (8″) we found the best fit for 211 Å has $\langle Q\rangle\approx 7\times 10^{-3}(B_{\mbox{\scriptsize avg}}/{100\mbox{G}})^{1.5}(L/{10^{9}\mbox{cm}})^{-1}$ erg cm-3 s-1, and for 335 Å the best fit is $\langle Q\rangle\approx 4\times 10^{-3}(B_{\mbox{\scriptsize avg}}/{100\mbox{G}})^{0.5}(L/{10^{9}\mbox{cm}})^{-1.5}$ erg cm-3 s-1. However, for all fits there is a range of parameter values that give qualitatively reasonable fits lying roughly in a line in ($a,b$) space, indicative of a dependence between $L$ and $B_{\mbox{\scriptsize avg}}$. This line was approximately defined by $b\approx 1.6-0.6a$. This relationship is consistent with a $B_{\mbox{\scriptsize avg}}$ dependence on $L$.
@@
We also found that for EUV lines showing a cooler (but still optically thin) plasma the models did not replicate the emission from loops along the edge of the active region, which dominates the emission in the 171 and 131 Å bands and also contributes significantly to the 94 and 193 Å bands. We conclude that this emission is due to relatively high-altitude transition-region emission in these long loops. EBTEL defines the transition region to be the section of the loop in which thermal conduction serves to heat the plasma rather than cool it (as occurs in the coronal section), and as such the transition region generally fills a length about 5% of the full loop length from the footpoint. GX Simulator and the EBTEL model informing it assume that all transition-region emission comes from one voxel at the footpoint, which is not sufficient to model the spatial distribution of the emission in these AIA bands. Thus we conclude that it is important in modeling active regions to consider the spatial extent of the upper transition region as a function of temperature. The cooler part of the transition region is confined to low altitudes close to the footpoints.
@@
T.A.K. and J.A.K. were supported by the GSFC Heliophysics Internal Scientist Funding Model (competitive work package program). This work is also part of the effort of the Committee on Space Research International Space Weather Action Team (COSPAR-ISWAT) S2-06: Origins of the Spectral Irradiance and its Intermediate Timescale Variability. GX Simulator is a package available as part of Solar Soft (Freeland and Handy, 1998) and also on github at https://github.com/Gelu-Nita/GX_SIMULATOR. G.D.F. and G.M.N. were partly supported by the NSF grants RISE-2324724 and AST-2206424, NASA grant 80NSSC23K0090 to the New Jersey Institute of Technology and by CSTR/NJIT. The authors gratefully acknowledge the use of data from SDO/AIA and HMI. Thanks to the referee for thoughtful suggestions that have significantly enhanced the paper. SDO (AIA, HMI)
Appendix A Transition-Region Mask
We present here a short physical justification for the transition-region mask (see Section II.5), which is expanded upon in more detail by Nita et al. (2018) in their Section 4.1 and Appendix A.
For a given loop length and heating rate, the brightness density of the transition-region emission at the solar surface decreases as the loop leg becomes more inclined to vertical, as discussed after Equation 5 of Nita et al. (2018). One way to understand this is as follows. The radiative energy losses from the transition region per unit cross-sectional area depend on the loop length and heating rate and are independent of the base inclination. By definition, the cross section is perpendicular to the loop axis. It is different from the footprint of the loop on the solar surface, which is larger than the cross section by an amount that increases with the tilt from vertical. Because the fixed radiative losses are spread out over a larger area, the brightness density of the footprint decreases with increasing tilt.
Photospheric magnetograms of active regions reveal regions of strong vertical magnetic field surrounded by regions of much weaker vertical field (e.g., Klimchuk, 1987). This gives rise to a magnetic canopy, where the strong vertical fields bend over sharply to become nearly horizontal in the chromosphere and transition region above the weak fields. From the discussion above, the transition region is observed to be much fainter above weak fields than strong fields.
Canopies do not arise in extrapolated magnetic field models because the transition from large to small plasma beta happens discontinuously at the lower boundary of the model. The field is much less inclined above regions of photospheric weak field than on the real Sun and therefore the predicted transition-region emission is much brighter. To correct this, we impose a transition-region mask. We set the transition-region emission to zero wherever the photospheric field is weaker than a critical value, which we take to be 200 G.
Appendix B AIA Model Grids
@@
This appendix contains more examples of calculated models that give the lowest $\sum\delta^{2}$ values as a function of $q_{0}$ for each of a grid of ($a,b$) values, Figures 18-21.
@@
Figure 18: Thumbnails of model 211 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for calculations using Magnetic Model 1, a model/data comparison resolution of 8″, and including both the corona and transitions region. The actual 211 Å image at the same resolution is on the bottom left. The $a=1.5$, $b=1$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. See Fig. 8b for larger version of the best model. Maps with $\sum\delta^{2}$ within 10% and 20% of the minimum have titles and borders in orange and red, respectively. The images are all scaled linearly with the same intensity range.
@@
Figure 19: Thumbnails of model 211 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for calculations using Magnetic Model 1, a model/data comparison resolution of 2″, and including the corona only. The actual 211 Å image at the same resolution is at the bottom left. The $a=2$, $b=0.5$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. See Fig. 9b for larger version of the best model. Maps with $\sum\delta^{2}$ within 10% and 20% of the minimum have titles and borders in orange and red, respectively. The images are all scaled linearly with the same intensity range.
@@
Figure 20: Thumbnails of model 171 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for calculations using Magnetic Model 2, a model/data comparison resolution of 2″, and including the corona and transition region. The actual 171 Å image at the same resolution is at the bottom left. The $a=1.0$, $b=0$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. See Fig. 13b for larger version of the best model. Maps with $\sum\delta^{2}$ within 10% and 20% of the minimum have titles and borders in orange and red, respectively. The images are all scaled linearly with the same intensity range.
Figure 21: Thumbnails of model 335 Å images showing the best $q_{0}$ solutions for each of the ($a,b$) pairs considered for calculations using Magnetic Model 1, a model/data comparison resolution of 8″, and including both the corona and transitions region. The actual 335 Å image at the same resolution is on the bottom left. The $a=0.5$, $b=1.5$ image has the lowest $\sum\delta^{2}$ value. Its title and border are in green. See Fig. 14 bottom right panel for larger version of the best model. Maps with $\sum\delta^{2}$ within 10% and 20% of the minimum have titles and borders in orange and red, respectively. The images are all scaled linearly with the same intensity range.